# Qingchuan Yang 1,  Simon Mahns  [^2] footnotemark: 2,  Sida Li 322footnotemark: 2,  Anri Gu 322footnotemark: 2,  Jibang Wu 4,  Haifeng Xu  [^3] footnotemark: 3

### v1

@@ 1 Introduction
Forecasting is a fundamental intellectual pursuit that has shaped human progress from the earliest scientific inquiries to modern economics and finance. In machine learning, forecasting has also been a central theme, with rich traditions ranging from time-series analysis (Box & Jenkins, 1976; Lim & Zohren, 2021), online learning (Foster & Vohra, 1999), to conformal prediction (Barber et al., 2023). Yet, somewhat surprisingly, the challenge of open-domain forecasting, producing accurate predictions across a wide range of topics without domain-specific tuning or specialized datasets, remains largely unexplored. Achieving reliable foresight in this setting would represent a qualitative leap in AI capability, with far-reaching societal implications, from enhancing market efficiency to guiding high-stakes policy decisions (Arrow et al., 2008).
@@
At its core, forecasting is the process of connecting present knowledge to anticipate future outcomes. Large language models (LLMs) seem natural candidates for this role. Trained on massive corpora of human knowledge through the seemingly narrow objective of next-word prediction, LLMs have developed emergent capabilities that extend far beyond their training objective (Bubeck et al., 2023). This motivates the hypothesis that the ability to predict the next word may also give rise to the ability to predict the next event. If substantiated, this would position LLMs not only as repositories of human knowledge but also as instruments of reliable foresight, leading to the prospect of LLM-as-a-Prophet: Can AI systems reliably predict the future by connecting the dots across existing real-world information?
@@
In this paper, we seek to systematically examine the prospects and challenges of building general-purpose systems for open-domain forecasting. On one hand, forecasting is a natural next pursuit given the rapid progress of AI, as it draws on a combination of advanced capabilities that current models are only beginning to demonstrate: information retrieval, complex reasoning and data analysis. Moreover, as many established benchmarks are approaching saturation and are increasingly prone to training-data contamination (Deng et al., 2024), open-domain forecasting provides a forward-looking and contamination-free setting with objectively measurable outcomes, making it a rigorous testbed for evaluating advanced model intelligence (Zou et al., 2022; Karger et al., 2025). On the other hand, we observe that current LLMs often struggle with key requirements for reliable foresight, including calibrated uncertainty estimation (Geng et al., 2023) and robust reasoning (Zhou et al., 2024) in the presence of noisy or incomplete evidence. As a result, their forecasting results may at times resemble guesswork rather than deliberated prediction, raising the possibility that fundamental barriers must be addressed before such evaluations can serve as a meaningful benchmark at the present time (Paleka et al., 2025a). Toward this end, we introduce Prophet Arena, a general framework for evaluating LLMs on live, real-world forecasting questions in a controlled and extensible way. Our goal is not only to assess the current forecasting performance of LLMs, but also to use forecasting as a lens for studying core components of intelligence, including reasoning, calibration, evidence aggregation. By doing so, we aim to identify which capabilities are emerging, which remain limited, and how forecasting evaluation can guide the development of more reliable predictive intelligence.
@@
Understanding LLMs’ Forecasting Capabilities. The central goal of our work is to understand the novel paradigm of LLM-as-a-Prophet and analyze how different capabilities (e.g., knowledge internalization, source usage, etc.) affect LLMs’ predictive intelligence. This research goal shares similarity to a few recent works on understanding and diagnosing special aspects of LLMs’ forecasting capabilities. For instance, (Dai et al., 2025; Zhu et al., 2025) study LLMs’ temporal generalization capability by challenging LLMs to forecast future events curated from news articles. Both works show that LLMs’ forecasting accuracy degrades over time, even when they are armed with retrieval augmented generation (RAG). Paleka et al. (2025b) study whether LLMs can make consistent forecasts. For example, a logical AI should not predict that both the Democratic and Republican parties have 60% probability of winning the 2024 US presidential election. Towards that end, they build a proper-scoring-rule forecasting benchmark to measure the consistency of LLMs’ predictions. Similar to both works’ research methodology, we also build a benchmark platform Prophet Arena to study our research question. However, to our knowledge, our work is the first to investigate the general predictive intelligence of the LLM-as-a-Prophet paradigm.
@@
Forecasting Benchmarks. Forecasting has recently become a popular challenge for benchmarking LLMs’ capabilities. Besides being a real challenge to LLM, it also avoids the thorny issue of benchmark contamination due to the evaluating LLMs on “future” events (Dai et al., 2025; Karger et al., 2025). To our knowledge, (Jin et al., 2021) is perhaps one of the earliest to test such forecasting capabilities of language models. They introduced ForecastQA, an evaluation dataset consisting of 10,392 crowdsourced multiple-choice questions, and the performance of language models at the time (mainly BERT models) still significantly lags behind human performance. Since (Jin et al., 2021), there has been a progressive line of work developing more and more challenging forecasting benchmarks for more advanced models, by integrating prediction market events (Zou et al., 2022), extracting events from news articles (Zhang et al., 2024; Wang et al., 2025), curating future-oriented questions from websites (Wildman et al., 2025), using open-ended queries (Guan et al., 2024), and lately developing dynamic benchmarks with live event and leaderboard updates (Zeng et al., 2025; Karger et al., 2025; Bianchi et al., 2025).
We also build a benchmark Prophet Arena, but as mentioned above, the main goal of our work is to understand LLM-as-a-prophet and analyze how its different capabilities affect the forecasting. We built Prophet Arena mainly for this analysis purpose, similar to that of (Dai et al., 2025; Paleka et al., 2025b). Towards this end, our work departs from these benchmark researches in multiple ways. First, we focus on comprehensive evaluation about LLMs’ predictive intelligence from various dimensions such as the quality, trustworthiness and economic values of forecasts. In contrast, the works above mainly focus on curating the right dataset but do not dive deep into evaluation metrics. In fact, almost all the aforementioned works only report forecasting accuracy, which cannot fully capture the quality of forecasting (as we illustrate in Section 3). The only exception is (Karger et al., 2025), which uses Brier scores (one of the three metrics we examine). Second, because our goal is not to rank LLMs or agents as in the above benchmark papers but rather to understand frontier LLMs’ forecasting capabilities, we need a natural baseline forecaster to know how good frontier LLMs are. Towards that end, we build Prophet Arena based on events from prediction markets, and hence can use the market’s forecasts as a cost-effective and justified baseline forecaster (Arrow et al., 2008). Finally, besides reporting different evaluation metrics, we further dive into the LLM-as-a-Prophet paradigm and offer insights about how the forecasts are affected by different capabilities such as knowledge internalization and source usage.
@@ 2.3 Design Choices of  Prophet Arena and Differences from Other Forecasting Benchmarks
Because our goal is to thoroughly evaluate what capabilities affect LLMs’ predictive intelligence, Prophet Arena is built with different design choices compared to several recent forecasting benchmarks, which are designed primarily to rank LLMs or agents based solely on forecasting accuracy. We further illustrate these differences in Table 1, whereas highlight the four key design choices of Prophet Arena below.
@@
    Probabilistic forecasts: Future events are intrinsically random, so we elicit each market’s probabilistic forecast from each model, rather than a single choice of the most likely outcome. Notably, probabilistic forecasting is also the standard for real-world forecasting platforms in general, including leading ones such as Metaculus (2015) and Good-Judgement-Open (2015).
@@
    Multi-horizon protocol: For each event, the models are asked to make prediction across various timestamps before the event resolves. This enables temporal analyses of information assimilation and model behaviors.
@@
Table 1: Comparison between Prophet Arena and related forecasting benchmarks. We compare against MIRAI (Ye et al., 2024), ForecastBench (Karger et al., 2025), FutureBench (Bianchi et al., 2025), and FutureX (Zeng et al., 2025).
@@ 3 Evaluating Forecasts: Caveats, Metrics, and Results
The golden metric in machine learning for evaluating prediction of binary (or multiary) outcomes is the accuracy measure (i.e., 0-1 loss). Indeed, several recent studies on benchmarking LLMs’ forecasting capabilities (e.g., Zeng et al. (2025); Karger et al. (2025); Wildman et al. (2025)) have used models’ accuracy as an indicator of their forecasting capabilities. These studies ask LLMs to predict a deterministic outcome and then evaluate this prediction’s accuracy (i.e., 0-1 loss). As a first caveat, we point out the limitation of measuring probabilistic forecasts via the accuracy metric. This limitation is due to the random nature of future events, which make them intrinsically different from typical classification tasks (e.g., recognizing whether a cat is in an image or not) with a deterministic answer. This is also why real-world forecasting platforms, including leading ones such as Metaculus (2015) and Good-Judgement-Open (2015), almost always elicit probabilistic forecasting. [^3] Some classification algorithms, such as logistic regression, also compute probabilities. However, unlike forecasting, such probabilities are not a requirement on the output but rather for the sake of algorithm design. Indeed, these algorithms often convert the probability to a deterministic output as its final prediction.
Caveat 1: Accuracy cannot fully measure the quality of probabilistic forecasts. To see this, consider a binary random event $E$ with groundtruth probability $p^{*}=0.6$. Suppose Alice has a perfect forecast of the probability as $p^{A}=0.6$ whereas Bob’s forecast is $p^{B}=1$. To evaluate their predictions’ accuracy, each forecaster must give a predicted outcome $o\in\{0,1\}$. In this case, it is natural for both to predict $o=1$ (as they both believe $o=1$ is more probable), which yields the same accuracy $\mathbb{E}_{o\sim p^{*}}\Pr(o=1)=0.6$. Therefore, the accuracy metric fails to distinguish Alice’s perfect forecast of the event’s probability from Bob’s forecast, which is extremal and not precise. The intrinsic reason of this failure is that using accuracy alone ignores prediction confidence, which is reflected in forecasted probabilities. In this example, Bob is overly confident about the outcome $o=1$ compared to Alice, despite having the same accuracy as her. Nevertheless, correctly measuring the confidence of a forecast is crucial for downstream decision making tasks, such as risk control in financial markets (Rigotti & Shannon, 2005).
@@
Given the Caveat 1 above, we identify three different dimensions of forecasting evaluation in order to obtain a comprehensive understanding of LLMs’ forecasting capabilities. These three dimensions are (1) forecasting loss (Gneiting & Raftery, 2007; Brier, 1950) which measures the absolute quality of a probabilistic prediction; (2) calibration error (DeGroot & Fienberg, 1983; Guo et al., 2017) which measures trustworthiness (formally, statistical consistency) of a probabilistic prediction; and (3) market return (Mallikarjuna & Rao, 2019) which measures the relative advantage over current market’s consensus. [^4] Market consensus about an event’s forecast is generally difficult to obtain. However, a good proxy of such data is available for Prophet Arena as it fetches forecasting events from prediction markets, which are widely believed to offer good approximation of the market’s consensus on the event’s probability (Arrow et al., 2008; Berg et al., 2008). As we illustrate below, these dimensions capture fundamentally different aspects of a forecast, enabling us to build a comprehensive understanding of LLMs predictive intelligence. For each dimension, we choose the most standard metrics.
@@ 3.1.1 Scoring Rules to Measure Forecasting Loss
The standard approach for evaluating probabilistic forecasts is through proper scoring rules (Gneiting & Raftery, 2007), which quantifies the discrepancy between predicted probabilities and realized outcomes. A scoring rule is proper if it ensures that the perfect probabilistic forecast achieves the optimal expected score. Prophet Arena adopts one of the most popular proper scoring rules, the Brier score (Brier, 1950), which is defined for each event $E_{i}$ as [^5] Careful readers might notice that definition given here has some subtle difference from a “textbook version.” We delay a detailed explanation to § B.1.
@@
3.1.2 Calibration Errors to Measure Trustworthiness
In addition to achieving low Brier scores, a good forecast should also be trustworthy – that is, when it assigns probability $\tilde{p}$ to Yes, the event should indeed occur about $\tilde{p}$ fraction of the time. This notion of trustworthiness is conventionally studied in the machine learning literature as calibration (DeGroot & Fienberg, 1983; Guo et al., 2017; Kalai & Vempala, 2024).
@@ 3.1.2 Calibration Errors to Measure Trustworthiness
Intuitively, ECE captures how much a probabilistic forecast differs from the real averaged probability, given this forecast level. A lower ECE means the forecast is more trustable, though it does not necessarily imply small forecasting loss. This difference is reflected in our LLM evaluations below, whereas an illustrative mathematical example can be found in § B.2.

### v2

@@ 1 Introduction
Forecasting is a fundamental intellectual pursuit that has shaped human progress from the earliest scientific inquiries to modern economics and finance. In machine learning, forecasting has also been a central theme, with rich traditions ranging from time-series analysis (box1976time; lim2021time), online learning (foster1999regret), to conformal prediction (barber2023conformal). Yet, somewhat surprisingly, the challenge of open-domain forecasting, producing accurate predictions across a wide range of topics without domain-specific tuning or specialized datasets, remains largely unexplored. Achieving reliable foresight in this setting would represent a qualitative leap in AI capability, with far-reaching societal implications, from enhancing market efficiency to guiding high-stakes policy decisions (arrow2008promise).
@@
At its core, forecasting is the process of connecting present knowledge to anticipate future outcomes. Large language models (LLMs) seem natural candidates for this role. Trained on massive corpora of human knowledge through the seemingly narrow objective of next-word prediction, LLMs have developed emergent capabilities that extend far beyond their training objective (bubeck2023sparks). This motivates the prospect that the ability to predict the next word may also give rise to the ability to predict the next event. Indeed, recent works by zou2022forecasting; halawi2024approaching have already made promising progress in employing language models for forecasting tasks. We envision that, with growing interest and continued progress, this would position LLMs not only as repositories of human knowledge but also as instruments of reliable foresight, leading to the prospect of LLM-as-a-Prophet: Can AI systems reliably predict the future by connecting the dots across existing real-world information?
@@
In this paper, we seek to systematically examine the prospects and challenges of building general-purpose systems for open-domain forecasting. On one hand, forecasting is a natural next pursuit given the rapid progress of AI, as it draws on a combination of advanced capabilities that current models are only beginning to demonstrate: information retrieval, complex reasoning and data analysis. Moreover, as many established benchmarks are approaching saturation and are increasingly prone to training-data contamination (deng2024investigating), open-domain forecasting provides a forward-looking and contamination-free setting with objectively measurable outcomes, making it a rigorous testbed for evaluating advanced model intelligence (zou2022forecasting; karger2024forecastbench). On the other hand, we observe that current LLMs often struggle with key requirements for reliable foresight, including calibrated uncertainty estimation (geng2023survey) and robust reasoning (zhou2024can) in the presence of noisy or incomplete evidence. As a result, their forecasting results may at times resemble guesswork rather than deliberated prediction, raising the possibility that fundamental barriers must be addressed before such evaluations can serve as a meaningful benchmark at the present time (paleka2025evaluating). Toward this end, we introduce Prophet Arena, a general framework for evaluating LLMs on live, real-world forecasting questions in a controlled and extensible way. Our goal is not only to assess the current forecasting performance of LLMs, but also to use forecasting as a lens for studying core components of intelligence, including reasoning, calibration, evidence aggregation. By doing so, we aim to identify which capabilities are emerging, which remain limited, and how forecasting evaluation can guide the development of more reliable predictive intelligence.
@@
Understanding and advancing LLMs’ forecasting capabilities. A key goal of our work is to understand the novel paradigm of LLM-as-a-Prophet and analyze how different capabilities (e.g., knowledge internalization, source usage, etc.) affect LLMs’ predictive intelligence. This research goal shares similarities to a few recent works on understanding and diagnosing special aspects of LLMs’ forecasting capabilities. For instance, daillms; chenghaozhu2025your study LLMs’ temporal generalization capability by challenging LLMs to forecast future events curated from news articles. Both works show that LLMs’ forecasting accuracy degrades over time, even when they are armed with retrieval augmented generation (RAG). paleka2025consistency studies whether LLMs can make consistent forecasts; for example, a logical AI should not predict that both the Democratic and Republican parties have a 60% chance of winning the 2024 US presidential election. Towards that end, they build a proper-scoring-rule forecasting benchmark to measure the consistency of LLMs’ predictions. Similar to both works’ research methodology, we also build a benchmark platform Prophet Arena to study our research question. However, to our knowledge, our work is the first to investigate general predictive intelligence within the LLM-as-a-Prophet paradigm. Finally, we also note the recent line of works advancing the forecasting capabilities of AI systems using language models (zou2022forecasting; halawi2024approaching). Though our goal of benchmarking is different, we envision that insights from our analysis about the strengths and limitations of current LLMs for forecasting could benefit future research for advancing AI’s forecasting capabilities.
@@
Forecasting Benchmarks. Forecasting has recently become a popular challenge for benchmarking LLMs’ capabilities. Besides being a real challenge to LLMs, it also avoids the thorny issue of benchmark contamination due to the evaluating LLMs on “future” events (daillms; karger2024forecastbench). To our knowledge, (jin2021forecastqa) is perhaps one of the earliest to test such forecasting capabilities of language models. They introduced ForecastQA, an evaluation dataset consisting of 10,392 crowdsourced multiple-choice questions, and the performance of language models at the time (mainly BERT models) still significantly lags behind human performance. Since jin2021forecastqa, there has been a progressive line of work developing more and more challenging forecasting benchmarks for more advanced models, by integrating prediction market events (zou2022forecasting), extracting events from news articles (zhang2024analyzing; wang2025openforecast), curating future-oriented questions from websites (wildman2025bench), using open-ended queries (guan2024openep), and lately developing dynamic benchmarks with live event and leaderboard updates (zeng2025futurex; karger2024forecastbench; FutureBench).
We also develop a benchmark, Prophet Arena, which we intentionally ground on thousands of prediction market events, leveraging their incentive-aligned participation and standardized resolution criteria to obtain a more rigorous testbed that mitigates selection bias from question design. However, our goal goes beyond ranking models: we aim to understand the LLM-as-a-Prophet paradigm and analyze how different model capabilities affect forecasting ability. We build Prophet Arena primarily for this analytical purpose, similar to daillms; paleka2025consistency.
Towards this end, our work departs from previous benchmarks in several ways. First, we provide a comprehensive evaluation of LLMs’ predictive intelligence across multiple dimensions, including Brier score, calibration, and economic value. In Section 3.1.1, we illustrate how these metrics differ and when each should be preferred. In contrast, most prior work primarily focuses on a single metric, such as Brier score (halawi2024approaching; karger2024forecastbench), calibration (zou2022forecasting), or accuracy (zeng2025futurex; daillms), which does not fully capture forecast quality (as we further demonstrate in Section 3). Second, as our goal is to better understand frontier LLMs’ forecasting capabilities, we require a natural baseline forecaster. By grounding Prophet Arena in prediction market events, we can use market-implied probabilities as a cost-effective and theoretically justified baseline (arrow2008promise). Finally, beyond reporting evaluation metrics, we further analyze the LLM-as-a-Prophet paradigm and provide insights into how forecasts are shaped by specific capabilities such as knowledge internalization and source usage.
@@ 2.3 Design Choices of  Prophet Arena and Differences from Other Forecasting Benchmarks
Because our goal is to thoroughly evaluate what capabilities affect LLMs’ predictive intelligence, Prophet Arena is built with different design choices compared to several recent forecasting benchmarks, which are designed primarily to rank LLMs or agents based solely on forecasting accuracy. To our best knowledge, Prophet Arena is the first large-scale benchmark that continuously runs evaluation on real-market events under a multi-horizon protocol and within a modularized forecasting pipeline. We further illustrate these differences in Table 1, whereas highlight the four key design choices of Prophet Arena below.
@@
    Probabilistic forecasts: Future events are intrinsically random, so we elicit each market’s probabilistic forecast from each model, rather than a single choice of the most likely outcome. Notably, probabilistic forecasting is also the standard for real-world forecasting platforms in general, including leading ones such as Metaculus and GoodJudge.
@@
    Multi-horizon protocol: We implement a multi-horizon forecasting protocol that sets a schedule for the models to make prediction across various timestamps before the event resolves; we defer details on the scheduling algorithm and multi-horizon aggregation to § A.2. This enables our temporal analyses of how models update their forecasts as market conditions and public information evolve across the full lifecycle of an event.
@@
Table 1: Comparison between Prophet Arena and related forecasting benchmarks. We compare against MIRAI (DBLP:journals/corr/abs-2407-01231), ForecastBench (karger2024forecastbench), FutureBench (FutureBench), and FutureX (zeng2025futurex).
@@ 3 Evaluating Forecasts: Caveats, Metrics, and Results
The golden metric in machine learning for evaluating prediction of binary (or multiary) outcomes is the accuracy measure (i.e., 0-1 loss). Indeed, several recent studies on benchmarking LLMs’ forecasting capabilities (e.g., zeng2025futurex; karger2024forecastbench; wildman2025bench) have used models’ accuracy as an indicator of their forecasting capabilities. These studies ask LLMs to predict a deterministic outcome and then evaluate this prediction’s accuracy (i.e., 0-1 loss). As a first caveat, we point out the limitation of measuring probabilistic forecasts via the accuracy metric. This limitation is due to the random nature of future events, which make them intrinsically different from typical classification tasks (e.g., recognizing whether a cat is in an image or not) with a deterministic answer. This is also why real-world forecasting platforms, including leading ones such as Metaculus and GoodJudge, almost always elicit probabilistic forecasting. [^3] Some classification algorithms, such as logistic regression, also compute probabilities. However, unlike forecasting, such probabilities are not a requirement on the output but rather for the sake of algorithm design. Indeed, these algorithms often convert the probability to a deterministic output as its final prediction.
Caveat 1: Accuracy cannot fully measure the quality of probabilistic forecasts. To see this, consider a binary random event $E$ with groundtruth probability $p^{*}=0.6$. Suppose Alice has a perfect forecast of the probability as $p^{A}=0.6$ whereas Bob’s forecast is $p^{B}=1$. To evaluate their predictions’ accuracy, each forecaster must give a predicted outcome $o\in\{0,1\}$. In this case, it is natural for both to predict $o=1$ (as they both believe $o=1$ is more probable), which yields the same accuracy $\mathbb{E}_{o\sim p^{*}}\Pr(o=1)=0.6$. Therefore, the accuracy metric fails to distinguish Alice’s perfect forecast of the event’s probability from Bob’s forecast, which is extremal and not precise. The intrinsic reason of this failure is that using accuracy alone ignores prediction confidence, which is reflected in forecasted probabilities. In this example, Bob is overly confident about the outcome $o=1$ compared to Alice, despite having the same accuracy as her. Nevertheless, correctly measuring the confidence of a forecast is crucial for downstream decision making tasks, such as risk control in financial markets (rigotti2005uncertainty).
@@
Given the Caveat 1 above, we identify three different dimensions of forecasting evaluation in order to obtain a comprehensive understanding of LLMs’ forecasting capabilities. These three dimensions are (1) forecasting loss (gneiting2007strictly; glenn1950verification) which measures the absolute quality of a probabilistic prediction; (2) calibration error (degroot1983comparison; guo2017calibration) which measures reliability (formally, statistical consistency) of a probabilistic prediction; and (3) market return (mallikarjuna2019evaluation) which measures the relative advantage over current market’s consensus. [^4] Market consensus about an event’s forecast is generally difficult to obtain. However, a good proxy of such data is available for Prophet Arena as it fetches forecasting events from prediction markets, which are widely believed to offer good approximation of the market’s consensus on the event’s probability (arrow2008promise; berg2008results). As we illustrate below, these dimensions capture fundamentally different aspects of a forecast, enabling us to build a comprehensive understanding of LLMs predictive intelligence. For each dimension, we choose the most standard metrics.
@@ 3.1.1 Scoring Rules to Measure Forecasting Loss
The standard approach for evaluating probabilistic forecasts is through proper scoring rules (gneiting2007strictly), which quantifies the discrepancy between predicted probabilities and realized outcomes. A scoring rule is proper if it ensures that the perfect probabilistic forecast achieves the optimal expected score. Prophet Arena adopts one of the most popular proper scoring rules, the Brier score (glenn1950verification), which is defined for each event $E_{i}$ as [^5] Careful readers might notice that definition given here has some subtle difference from a “textbook version.” We delay a detailed explanation to § B.1.
@@
3.1.2 Calibration Errors to Measure Reliability
In addition to achieving low Brier scores, a good forecast should also be reliable – that is, when it assigns probability $\tilde{p}$ to Yes, the event should indeed occur about $\tilde{p}$ fraction of the time. This notion of reliability of a forecaster dates back to (murphy1973new) and has been conventionally studied in the machine learning literature as calibration (degroot1983comparison; guo2017calibration; kalai2024calibrated)  [^6] We note that recent works on forecast evaluations have separately considered Brier scores (e.g., (halawi2024approaching; karger2024forecastbench)) and calibration errors (e.g., (zou2022forecasting)), though no work has conducted comprehensive evaluation of all these metrics to our knowledge, especially models’ market returns..
@@ 3.1.2 Calibration Errors to Measure Reliability
Intuitively, ECE captures how much a probabilistic forecast differs from the real averaged probability, given this forecast level. A lower ECE means the forecast is more reliable, though it does not necessarily imply small forecasting loss. This difference is reflected in our LLM evaluations below, whereas an illustrative mathematical example can be found in § B.3.